OpenStudy (xoxolove456):

Chris and Melissa can clean a room together in 5 hours. Working alone Chris can clean the room in 1 hour less time than Melissa can.

OpenStudy (xoxolove456):

How long does it take Melissa to clean the room alone ?

OpenStudy (xoxolove456):

@jim_thompson5910

OpenStudy (dumbcow):

Rate that they work is inverse of time. Together they work at a rate of 1/5 room per hour Let c,m be the respective times \[\frac{1}{c} + \frac{1}{m} = \frac{1}{5}\] \[c = m-1\] \[\rightarrow \frac{1}{m-1} + \frac{1}{m} = \frac{1}{5}\] solve for m \[\frac{m + (m-1)}{m(m-1)} = \frac{1}{5}\] \[2m - 1 = \frac{1}{5} (m^2 - m)\] \[m^2 - 11m +5 = 0\] \[m = \frac{11 +\sqrt{101}}{2}\]

OpenStudy (xoxolove456):

So how many hours ?

OpenStudy (dumbcow):

m hours